You are choosing between two health clubs. Club A offers a membership for a fee of $28 plus a monthly fee of $15.Club B offers a membership for a fee of $24 plus a monthly fee of $16.
After how many months will the total cost of each health club be the same? What will be the total cost for each club?
Please help. Thank you!

Answer by toidayma(44)

Let x be the number of months that for such period of time, the total cost for each club are equal.
We have:
the total cost for joining club A for x months is: 28 + 15x
the total cost for joining club B for x months is: 24 + 16x
Since the cost for joining each club for x month are equal, we have the equation:
28 + 15x = 24 + 16x <->
4 = x.
So, after 4 month since joining, the total cost of joining both club are the same. And that same total cost is 28 + 15*4 = 88 ($).